OpenStudy (ayee_ciera):

Find the square. (6m+7)^2

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle ({\rm \color{blue}{a}}+{\rm \color{red}{b}})^2={\rm \color{blue}{a}}^2+2{\rm \color{blue}{a}\color{red}{b}}+{\rm \color{red}{b}} }\) \(\large\color{slate}{ \displaystyle ({\rm \color{blue}{a}}-{\rm \color{red}{b}})^2={\rm \color{blue}{a}}^2-2{\rm \color{blue}{a}\color{red}{b}}+{\rm \color{red}{b}} }\)

OpenStudy (stamp):

~~~!! SPOILER!!! ANSWER BELOW !! ~~~ \[36m^2+84m+49\]

OpenStudy (ayee_ciera):

3(x^3 – 4x^2 + 9x) 3x(x^2 – 4x + 9) 3x(x^2 – 12x + 9) –3x(x^2 + 4x + 9)

OpenStudy (ayee_ciera):

answer choices^^^^^^